Three capacitors are connected in series: C1 = 28 μF, C 2 = 42 μF and C3 = 64 μF. What is the total capacitance?


1
Expert's answer
2022-02-25T11:11:54-0500

Answer

total capacitance

"C=\frac{ C_1* C_2*C_3}
 {C_2*C_3+C_1*C_3+C_1*C_2}\\=\frac{28*42*64*10^{-18}}{(28*42+42*64+64*28) *10^{-12}}\\=13.31*10^{-6}F"